A man borrows 850 taka from bank. The bill is renewed after every half year at an increase of 21% . After haw much time amount will become 8500 taka.


1
Expert's answer
2021-02-02T06:27:07-0500

Let the elapsed time be "n" years

For it renewed after every half year it will be "2n"


"A=8500"

"P=850"

"i=21"%

"A=p(1+i)"2n


"8500=850(1+21"%")"2n


"\frac{8500}{750}=(1+\frac{21}{100})"2n


Taking "log" on both sides,

"log(\frac{8500}{850}) =log(1+\frac{21}{100})"2n


"log(\frac{8500}{850}) =2nlog(1+\frac{21}{100})"


"log(\frac{8500}{850}) =2n log(\frac{121}{100})"


"1= 2n(0.08279)"


"n=\frac{1}{2×0.08279}"


"n=6.0393"


"n=6" years approximately